### v1

# Accretion of Generalized Chaplygin Gas onto Cosmologically Coupled Black Holes

Abstract
We study the accretion of cosmic dark fluids, responsible for driving the accelerated expansion of the universe, onto cosmologically coupled black holes. More specifically, we focus on the accretion of the Generalized Chaplygin Gas (GCG). To incorporate the global features of the GCG into this analysis, we employ the McVittie metric, which describes a black hole embedded in an expanding cosmological background. Within this framework, accretion is studied while consistently accounting for the backreaction on the metric components. Using a perturbative approach, we derive an expression for the effective black hole mass and for the evolution of both the black hole and cosmological apparent horizons under accretion. The analysis is performed in two distinct cosmological regimes: first, a matter-dominated era, and subsequently, a de Sitter era. In both cases, it is possible to determine analytically the instant in which accretion begins. For the matter-dominated era, the analytical expression shows that the greater the amount of matter available for accretion, the longer the accretion takes to start.
I Introduction
[…]
While the cosmological constant remains the simplest and most widely adopted model for dark energy, a variety of dynamical models have been proposed to address the theoretical and observational challenges. Among these, the Generalized Chaplygin Gas (GCG) has attracted attention due to its ability to interpolate between dust-like matter at early times and a dark energy–dominated regime at late times [kamenshchik2001alternative, Bilic:2001cg, bento2002generalized].
The GCG model not only provides a compelling observational framework to account for the universe’s accelerated expansion and to address the dark matter problem [bento2003wmap, aurich2018concordance, barreiro2008wmap], but also offers promising avenues for addressing fundamental issues such as the cosmological constant problem [bertolami2023seeding], the Hubble tension [yang2019dawn], the Swampland conjectures [bertolami2024sitter], and inflation through a model inspired on the GCG [Bertolami:2006zg].
Despite significant progress in constraining dark energy models, its interaction with compact astrophysical objects remains a relatively unexplored frontier. Black holes, with their intense gravitational fields and central role in galactic evolution, represent unique laboratories for probing such interactions under extreme conditions.
Some studies have already explored dark energy and the GCG in the context of compact objects, such as in Refs. [Mazur:2001fv, bertolami2005chaplygin, gorini2009more]. Other works have also investigated dark energy and the GCG in accretion scenarios, such as in Refs. [babichev2005accretion, rodrigues2012accretion], although there remains substantial room for further development in this area.
Classical accretion theory has been instrumental in describing the infall of baryonic matter and the associated electromagnetic signatures [bondi1952spherically, petterson1980variations, michel1972accretion, porth2017black]. However, in the context of a universe dominated by a dark energy, it becomes imperative to consider whether and how black holes may accrete this non-standard component.
The accretion of dark energy, particularly when modeled as a perfect fluid with negative pressure, introduces novel theoretical possibilities. In scenarios involving phantom energy, that satisfies the condition $p+\rho<-1$, it has been shown that the accretion process could lead to a decrease in black hole mass, a counterintuitive result that challenges conventional views of black hole growth and thermodynamics [babichev2004black]. However, despite the consistency of such approaches, certain aspects are often neglected, largely due to the considerable complexity involved in developing a comprehensive accretion model.
One such critical aspect is the backreaction induced by the accretion process itself. In the standard test-fluid approximation, where the accreting material is assumed too small to significantly modify the black hole’s spacetime and the geometry is treated as fixed, the accreting matter or energy is treated as having negligible influence on the background spacetime geometry.
However, for accretion processes that persist over cosmological timescales or involve significant energy fluxes, the assumption of a fixed black hole metric may no longer be valid. Backreaction effects can, in principle, modify the black hole’s horizon, influence the spacetime curvature in the near-horizon region, and potentially alter the global accretion dynamics. To this end, we employ a perturbative approach that enables to find solutions for the evolution of the black hole’s apparent horizon.
Another fundamental aspect is that, since accretion is inherently a local phenomenon, while dark energy models are employed to describe a global feature such as the accelerated expansion of the universe, it is essential to adopt a spacetime metric that accommodates both aspects. Specifically, we require a solution that locally describes a spherically symmetric object, while simultaneously incorporating the effects of cosmic expansion. In other words, we seek to study a black hole whose spacetime is not asymptotically flat, but rather asymptotically approaches a flat Friedmann-Lemaitre-Robertson-Walker (FLRW) cosmology.
For this purpose, we adopt the McVittie metric [mcvittie1933mass]. This spacetime solution has been the subject of ongoing debate for nearly eighty years regarding its physical consistency in describing black holes embedded in an expanding universe. However, recent studies have demonstrated that the McVittie metric not only provides a consistent framework for modeling black holes that are asymptotically FLRW [kaloper2010mcvittie, lake2011more, nandra2012effect, nandra2012effect1, faraoni2015cosmological], but also serves as an important foundational model for the development of more sophisticated and realistic metrics to this end [gao2011black].
The present work does not aim to delve into the broader debate concerning the validity and limitations of the McVittie metric. Nevertheless, since this metric will be employed in the next section to model both the accretion process and the associated backreaction, the results obtained may offer valuable insights and contribute to future efforts toward constructing more complete metrics capable of describing astrophysical objects within an expanding universe.
This paper is organized as follows: In Section II, We analyze the McVittie metric, with an emphasis on the effective mass associated with this geometry and on the time-dependent solutions of the apparent horizons. In Section III, we use a perturbative analysis that has already been developed that accounts for backreaction effects to study accretion within the McVittie spacetime. In Section IV, we apply the results developed in Section III to the GCG, focusing on two distinct analyses: one corresponding to a matter-dominated universe and the other to a de Sitter universe. Finally, in Section V, we present our conclusions and discuss the implications of the results.
Throughout this work we use units in which $c=G=1$ and adopt the metric signature to be +++-+++. Also, we will use the notation $^{\cdot}=\partial/\partial t$ and $^{\prime}=\partial/\partial r$.
II The McVittie metric and the GCG
The McVittie metric can be expressed in several equivalent forms; however, in this work, we adopt the formulation presented in Ref. [kaloper2010mcvittie]. The line element is given by
$ds^{2}=-\left(1-\frac{2M_{0}}{r}-H_{0}^{2}\left(t\right)r^{2}\right)dt^{2}-\frac{2rH_{0}\left(t\right)}{\sqrt{1-\frac{2M_{0}}{r}}}dtdr+\frac{dr^{2}}{1-\frac{2M_{0}}{r}}+r^{2}d\Omega_{2}^{2},$ (1)
where $M_{0}$ is the mass parameter associated with the central black hole, $H_{0}\left(t\right)\equiv\dot{a}(t)/a(t)$ is the Hubble parameter, $a(t)$ is the scale factor of the universe, and $d\Omega_{2}^{2}=d\theta^{2}+\sin^{2}\theta d\varphi^{2}$ is the metric element on the $2$-sphere. It is straightforward to verify that, in the limit $M_{0}\rightarrow 0$ and $r\rightarrow\infty$, the metric in Eq. (1) reduces to the FLRW metric [^1] In this work, we consider the McVittie metric in the context of a spatially flat universe., while for $H_{0}(t)\rightarrow 0$, it reduces to the Schwarzschild solution. Additionally, it is important to note that when $H_{0}(t)$ is a constant, the metric (1) corresponds to the Schwarzschild–de Sitter (SdS) solution.
Since we are not yet accounting for accretion, Eq. (1) will be treated as a background metric, denoted by $g_{\mu\nu}^{\left(0\right)}$, where the upper index ’$(0)$’ indicates the zeroth-order approximation in a perturbative expansion. Namely, the order that excludes accretion effects.
The Einstein field equations for metric (1) yield the following relations:
$-3H_{0}^{2}(t)$ $=$ $8\pi\left(T_{0}^{\text{ }0}\right)^{\left(0\right)},$ (2a)
$-3H_{0}^{2}-\frac{2\dot{H}_{0}}{\sqrt{1-\frac{2M_{0}}{r}}}$ $=$ $8\pi\left(T_{1}^{\text{ }1}\right)^{\left(0\right)}.$ (2b)
If we define $\left(T_{\nu}^{\text{ }\mu}\right)^{\left(0\right)}$ such that $\left(T_{0}^{\text{ }0}\right)^{\left(0\right)}=-\rho^{\left(0\right)}$ and $\left(T_{1}^{\text{ }1}\right)^{\left(0\right)}=p^{\left(0\right)}$, where $\rho^{\left(0\right)}$ and $p^{\left(0\right)}$ are the energy density and isotropic pressure, respectively, Eq. (2a) can be readily recognized as the Friedmann equation. This indicates that the presence of a black hole does not affect the homogeneity of the energy density distribution. However, Eq. (2b) reveals that the same is not true for the pressure, which acquires a spatial dependence due to the black hole’s influence.
To further elucidate the non-homogeneous nature of the pressure, we consider the energy conservation equation derived from the Einstein field equations. This relation is given by
$\dot{\rho}^{\left(0\right)}+3H_{0}\left(p^{\left(0\right)}+\rho^{\left(0\right)}\right)\sqrt{1-\frac{2M_{0}}{r}}=0.$ (3)
Since both $\rho^{\left(0\right)}$ and $H_{0}$ are functions of time only, we can infer that the term multiplying $3H_{0}$ in Eq. (3) must also be purely time-dependent. Therefore, we can write:
$\left(p^{\left(0\right)}+\rho^{\left(0\right)}\right)\sqrt{1-\frac{2M_{0}}{r}}=f\left(t\right),$ (4)
where $f\left(t\right)$ is an arbitrary function of time. The authors in Refs. [nandra2012effect, nandra2012effect1] have shown that for a matter-dominated universe, it is appropriate to take $f\left(t\right)=\rho^{\left(0\right)}$, such that the pressure can be written as
$p^{\left(0\right)}=\frac{\rho^{\left(0\right)}}{\sqrt{1-\frac{2M_{0}}{r}}}-\rho^{\left(0\right)}.$ (5)
In the limit $r\rightarrow\infty$, the pressure in Eq. (5) asymptotically approaches the form characteristic of a dust-like fluid.
We can also analyze Eq. (4) for a universe governed by the GCG. This fluid is described by an exotic equation of state of the form
$p_{GCG}\left(t\right)=-\frac{A}{\rho_{GCG}^{\alpha}\left(t\right)},$ (6)
where $A>0$ and $0<\alpha\leq 1$. The Chaplygin gas corresponds to the choice $\alpha=1$ [kamenshchik2001alternative, bento2002generalized, bento2003wmap]. In the framework of FLRW cosmology, the relativistic energy conservation associated to the equation of state (6) leads to an evolving energy density:
$\rho_{GCG}\left(t\right)=\left(A+\frac{B}{a^{3\left(1+\alpha\right)}\left(t\right)}\right)^{1/(1+\alpha)},$ (7)
where $B$ is an integration constant.
Therefore, in the study of a universe governed by the equation of state of the GCG, one is justified to choose in Eq. (4) $f\left(t\right)=\rho^{\left(0\right)}-A/\left(\rho^{\left(0\right)}\right)^{\alpha}$. With this identification, the pressure can be expressed as:
$p^{\left(0\right)}\left(r,t\right)=-\rho^{\left(0\right)}+\frac{1}{\sqrt{1-\frac{2M_{0}}{r}}}\left(\rho^{\left(0\right)}-\frac{A}{\left(\rho^{\left(0\right)}\right)^{\alpha}}\right),$ (8)
where far from the black hole, the pressure reduces to the standard GCG equation of state given in Eq. (6). Consequently, when the pressure is written in the form of Eq. (8), the solution to the energy conservation equation (3) recovers the same expression for $\rho^{\left(0\right)}$ as given in Eq. (7). However, for the sake of generality, we shall work with Einstein’s equations (2) in terms of the energy-momentum tensors components $\left(T_{\nu}^{\text{ }\mu}\right)^{\left(0\right)}$.
Other key aspects of the McVittie metric relevant to our analysis are the effective mass and the apparent horizon. The effective mass of the black hole can be computed using the Misner-Sharp mass function $M_{MS}^{\left(0\right)}$ [misner1964relativistic], defined by
$1-\frac{2M_{MS}^{\left(0\right)}}{r}=\left(g^{\mu\nu}\right)^{\left(0\right)}\partial_{\mu}r\partial_{\nu}r,$ (9)
which, using Eqs. (1) and (2a), results in
$M_{MS}^{\left(0\right)}\left(t,r\right)=M_{0}+\frac{H_{0}^{2}r^{3}}{2}.$ (10)
The apparent horizon of the black hole, denoted $r_{h}^{\left(0\right)}$, can be determined from the condition $g_{00}^{\left(0\right)}=0$, which, using Eq. (1), leads to the following cubic equation:
$H_{0}^{2}\left(r_{h}^{\left(0\right)}\right)^{3}-r_{h}^{\left(0\right)}+2M_{0}=0.$ (11)
As shown in Refs. [nickalls1993new, faraoni2012making] the solutions of the cubic equation above can be expressed as:
$r_{1}\left(t\right)$ $=$ $\frac{2}{\sqrt{3}H_{0}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{0}H_{0}\left(t\right)\right)}{3}\right),$ (12a)
$r_{2}\left(t\right)$ $=$ $\frac{1}{H_{0}\left(t\right)}\cos\left(\frac{\arcsin\left(3\sqrt{3}M_{0}H_{0}\left(t\right)\right)}{3}\right)$
$-\frac{1}{\sqrt{3}H_{0}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{0}H_{0}\left(t\right)\right)}{3}\right),$
$r_{3}\left(t\right)$ $=$ $-\frac{1}{H_{0}\left(t\right)}\cos\left(\frac{\arcsin\left(3\sqrt{3}M_{0}H_{0}\left(t\right)\right)}{3}\right)$
$-\frac{1}{\sqrt{3}H_{0}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{0}H_{0}\left(t\right)\right)}{3}\right).$
As discussed in Ref. [faraoni2012making], the solution $r_{3}$ can be discarded as it is unphysical, since it is always negative. The solution $r_{1}$ can be interpreted as the apparent horizon of the black hole, as it reduces to the Schwarzschild radius $r_{1}\rightarrow 2M_{0}$ in the limit $H_{0}\rightarrow 0$. Similarly, $r_{2}$ can be interpreted as the cosmological horizon, as it approaches $r_{2}\rightarrow 1/H_{0}$ when $M_{0}\rightarrow 0$. Henceforth we will denote $r_{1}$ by $r_{b}$ and $r_{2}$ by $r_{c}$.
Furthermore, the solutions given by Eqs. (12a) and (12) impose the following conditions: for $3\sqrt{3}M_{0}H_{0}<1$, the horiozons $r_{b}$ and $r_{c}$ are distinct; when $3\sqrt{3}M_{0}H_{0}=1$, the black hole horizon coincides with the cosmological horizon; and when $3\sqrt{3}M_{0}H_{0}>1$, no horizons exist and the black hole becomes a naked singularity.
Finally, it is also important to compute the expansion velocity of the apparent horizons. Differentiating (11) with respect to time yields the evolution velocity of the apparent horizon:
$\dot{r}_{i}=\frac{2H_{0}\dot{H}_{0}r_{i}^{3}}{1-3H_{0}^{2}r^{2}_{i}},$ (13)
where $i=b,c$.
This construction should be regarded as a simplified model, serving as a foundational framework for the development of more physically realistic approaches, particularly those that incorporate accretion processes. In the following section, we shall build upon the results presented thus far to construct a more refined model that accounts not only for accretion but also for the backreaction effects on the spacetime geometry.
III Perturbatitve approach to backreaction
The process of accretion within the framework of the McVittie metric has been investigated in previous works [gao2011black, gao2008does, guariento2012realistic]. The approach adopted in these studies primarily consists in promoting the mass parameter $M_{0}$ in Eq. (1) to a time-dependent function.
In the present work, we assume that the accretion of matter induces backreaction effects on the geometry. To investigate this backreaction, we employ a perturbative method as proposed in Ref. [babichev2012backreaction].
We begin by assuming that the backreaction induces both a temporal and radial dependence of the mass, as well as an inhomogeneity in the Hubble parameter. Consequently, the metric (1) becomes:
$ds^{2}=-\left(1-\frac{2M\left(t,r\right)}{r}-H\left(t,r\right)^{2}r^{2}\right)dt^{2}-\frac{2rH\left(t,r\right)}{\sqrt{1-\frac{2M\left(t,r\right)}{r}}}dtdr+\frac{dr^{2}}{1-\frac{2M\left(t,r\right)}{r}}+r^{2}d\Omega_{2}^{2}.$ (14)
The Einstein equations for the line element (14) then take the following form:
$8\pi\mathcal{T}_{0}^{\text{ }0}$ $=$ $-3H^{2}-2rHH^{\prime}-\frac{2M^{\prime}}{r^{2}},$ (15a)
$8\pi\mathcal{T}_{0}^{\text{ }1}$ $=$ $2rH\dot{H}+\frac{2\dot{M}}{r^{2}},$ (15b)
$8\pi\mathcal{T}_{1}^{\text{ }0}$ $=$ $-\frac{2\dot{M}}{\left(r-2M\right)^{2}},$ (15c)
$8\pi\mathcal{T}_{1}^{\text{ }1}$ $=$ $-\frac{1}{r^{2}\left(r-2M\right)^{2}}\left[3r^{2}H^{2}\left(r-2M\right)^{2}\right$.
$\left.+2\left(r-2M\right)\left(\left(r-2M\right)M^{\prime}+r^{3}\sqrt{1-\frac{2M}{r}}\dot{H}\right)\right$.
$\left.+2r^{3}H\left(\left(r-2M\right)^{2}H^{\prime}+\sqrt{1-\frac{2M}{r}}\dot{M}\right)\right].$
One also finds that the Einstein tensors $G_{2}^{\text{ }2}=G_{3}^{\text{ }3}\neq 0$; however, these components yield no independent equations, as they follow from Eqs. (15a)-(15c).
If we define the function
$\mathcal{M}_{MS}\left(t,r\right)=\frac{r}{2}\left(1-g_{00}\right)=M\left(t,r\right)+\frac{H\left(t,r\right)^{2}r^{3}}{2},$ (16)
where $g_{\mu\nu}$ is the metric associated with the line element (14), then Eqs. (15a) and (15b) can be rewritten as
$\mathcal{M}_{MS}^{\prime}\left(t,r\right)=\mathcal{B}\left(t,r\right)$ (17)
and
[…]
where
[…]
and
$\mathcal{B}\left(t,r\right)=-4\pi r^{2}T_{0}^{\text{ }0}\left(t,r\right).$ (20)
The system of partial differential equations (17)-(18) admits the general solution
$\mathcal{M}_{MS}\left(t,r\right)$ $=$ $M_{i}+\int_{t_{0ac}}^{t}\mathcal{A}\left(\bar{t},r_{0ac}\right)~d\bar{t}+\int_{r_{0ac}}^{r_{b}}\mathcal{B}\left(t,\bar{r}\right)~d\bar{r},$
where $t_{0ac}$ is chosen to be the instant in which accretion starts, $r_{0ac}$ is the apparent horizon at time $t=t_{0ac}$, $r_{b}$ is the geometric locus of the black hole apparent horizon, and $M_{i}$ is defined as $M_{i}=M_{MS}^{\left(0\right)}\left(t_{0ac},r_{0ac}\right)+M_{ac}\left(t_{0ac},r_{0ac}\right)$, with $M_{MS}^{\left(0\right)}\left(t,r\right)$ given by Eq. (10) and $M_{ac}\left(t_{0ac},r_{0ac}\right)$ being the mass of the accreted fluid at instant $t_{0ac}$ evaluated at $r_{0ac}$, such that $\mathcal{M}_{MS}\left(t_{0ac},r_{0ac}\right)=M_{i}$.
A solution similar to Eq. (LABEL:M_solution) appears in Refs. [babichev2012backreaction, campos2025dynamical]. However, in the present context, we are dealing with a slightly more general scenario, since both the accretion term and the energy density are allowed to depend on time. Moreover, it is worth highlighting that working with the McVittie metric allows us to obtain solutions analogous to that of Refs. [babichev2012backreaction, campos2025dynamical], but expressed in a physical time coordinate rather than in an Eddington–Finkelstein–type time coordinate.
The expression for $\mathcal{M}_{MS}$ reveals the contribution of the accretion of matter, described by $\mathcal{T}_{\nu}^{\text{ }\mu}$, to the initial effective mass $M_{i}$. The first integral in (LABEL:M_solution) corresponds to the contribution from the evolution of the accreting matter flux, evaluated in the region bounded by $r_{0ac}$ over the interval $t-t_{0ac}$. The second integral represents the time-dependent contribution from the energy density contained between the initial apparent horizon before accretion, $r_{0ac}$ , and the apparent horizon displaced due to accretion, $r_{b}$.
With the solution (LABEL:M_solution) at hand, we can now effectively compute the effects of accretion. We only need to evaluate the integrals in Eq.(LABEL:M_solution) in terms of the energy–momentum tensors. We shall assume that the tensors $\mathcal{T}_{\nu}^{\ \mu}$ satisfying Einstein’s equations (15) correspond to a perfect fluid, given by
$\mathcal{T}_{\nu}^{\ \mu}$ $=$ $\left({T}_{\nu}^{\ \mu}\right)^{(0)}+\left(\rho_{ac}+p_{ac}\right)u_{\nu}u^{\mu}+p_{ac}\delta_{\nu}^{\mu},$ (22)
where $\left({T}_{\nu}^{\ \mu}\right)^{(0)}$ represents the background fluid defined by equations (2), $\rho_{ac}$ and $p_{ac}$ are the energy density and the pressure of the accreting fluid, respectively, with $u^{\mu}$ its $4$-velocity.
We shall also assume that the accreting fluid has no angular motion, so that $u^{\mu}=\left(u^{0},vu^{0},0,0\right)$, where $v=dr/dt$. Using the normalization condition for the four-velocity, $u_{\mu}u^{\mu}=-1$, the relevant components of the energy–momentum tensor (22) are then given by
$\mathcal{T}_{0}^{\text{ }0}=-\frac{3H_{0}^{2}\left(t\right)}{8\pi}-\frac{\left(\rho_{ac}+p_{ac}\right)\left(g_{00}+g_{01}v\right)}{g_{00}+2g_{01}v+g_{11}v^{2}}+p_{ac}$ (23)
and
$\mathcal{T}_{0}^{\text{ }1}=-\frac{\left(\rho_{ac}+p_{ac}\right)\left(g_{00}v+g_{01}v^{2}\right)}{g_{00}+2g_{01}v+g_{11}v^{2}}.$ (24)
However, an evident difficulty immediately arises: the metric components are expressed in terms of the unknown functions $M\left(t,r\right)$ and $H\left(t,r\right)$, which appear in Eq. (14). To address this, we shall employ the perturbative approach used in Ref. [babichev2012backreaction]. We assume that the effects of the accreting fluid are much smaller than those of the background fluid, so that the energy–momentum tensor of the accreting fluid can be evaluated in the background metric $g_{\mu\nu}^{\left(0\right)}$ given by (1). Consequently, the tensor $\mathcal{T}_{\nu}^{\text{ }\mu}$ and the solution (LABEL:M_solution) will be considered as a first-order perturbation, namely
$\mathcal{T}_{\nu}^{\text{ }\mu}\rightarrow\left(T_{\nu}^{\text{ }\mu}\right)^{\left(1\right)},$ (25)
and
$\mathcal{M}_{MS}\rightarrow M_{MS}^{\left(1\right)}.$ (26)
For simplicity, we can analyze accretion at the apparent horizon of the black hole. At the black hole’s apparent horizon, $g_{00}^{\left(0\right)}=0$ and, consequently, from Eq. (1),
$1-\frac{2M_{0}}{r_{b}^{\left(0\right)}}=H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\right)^{2}.$ (27)
Eqs. (23) and (24) then reduce to:
$\left(T_{0}^{\text{ }0}\right)^{\left(1\right)}\left(t,r_{b}\right)$ $=$ $-\frac{3H_{0}^{2}\left(t\right)}{8\pi}+\frac{\left(\rho_{ac}+p_{ac}\right)H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}{v-2H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}$ (28)
$+p_{ac}$
and
$\left(T_{0}^{\text{ }1}\right)^{\left(1\right)}\left(t,r_{b}\right)=\frac{\left(\rho_{ac}+p_{ac}\right)H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}v}{v-2H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}.$ (29)
Under this perturbative approximation, Eq. (LABEL:M_solution) expressed in terms of the Eqs. (28) and (29) is exact at the initial time $t=t_{0ac}$, such that
$g_{00}\left(t_{0ac},r_{0ac}\right)=g_{00}^{\left(0\right)}\left(t_{0ac},r_{b}^{\left(0\right)}\left(t_{0ac}\right)\right),$ (30)
and provides an approximation for $t>t_{0ac}$.
We may then analyze the integrals appearing in Eq. (LABEL:M_solution) using the definitions from Eqs. (19) and (20). We are going to define the first integral in Eq. (LABEL:M_solution), after substituting Eq. (29), as
$F\left(t\right)\equiv\int_{t_{0ac}}^{t}\mathcal{A}\left(\bar{t},r_{0ac}\right)~d\bar{t}.$ (31)
Since our goal is to study the accretion of the GCG, before evaluating the second integral in Eq. (LABEL:M_solution), we anticipate that $\rho_{ac}$ and $p_{ac}$ with which we will work depend only on time. Evidently, the same assumption holds for the radial velocity of the fluid, $v$, once our goal is to evaluate the accretion at a specific point, the black hole horizon, i.e., $v(t,r)=v(t,r^{(1)}_{b})$. Consequently, $\left(T_{0}^{\text{ }0}\right)^{\left(1\right)}$ given by Eq. (28), will depend solely on time, and therefore Eq. (20) becomes
$\int_{r_{0ac}}^{r_{b}^{\left(1\right)}}\mathcal{B}\left(t,\bar{r}\right)~d\bar{r}$ $=$ $-4\pi\left(T_{0}^{\text{ }0}\left(t\right)\right)^{\left(1\right)}\int_{r_{0ac}}^{r_{b}^{\left(1\right)}}\bar{r}^{2}~d\bar{r}=$ (32)
$=$ $\frac{H_{1}^{2}\left(t\right)}{2}\left(r_{b}^{\left(1\right)}\right)^{3}+G\left(t\right),$
where we define
$H_{1}^{2}\left(t\right)\equiv-\frac{8\pi\left(T_{0}^{\text{ }0}\left(t\right)\right)^{\left(1\right)}}{3}$ (33)
and
$G\left(t\right)\equiv-\frac{H_{1}^{2}\left(t\right)}{2}r_{0ac}^{3}.$ (34)
Finally, the solution given in Eq. (LABEL:M_solution) can be expressed as
$M_{MS}^{\left(1\right)}\left(t,r\right)=M_{1}\left(t\right)+\frac{H_{1}^{2}\left(t\right)}{2}\left(r_{b}^{(1)}\right)^{3},$ (35)
where we define
$M_{1}\left(t\right)\equiv M_{i}+F\left(t\right)+G\left(t\right).$ (36)
Furthermore, this allows us to compute the evolution of the apparent horizon displaced due to accretion, $r^{\left(1\right)}$, through the equation $g_{00}^{\left(1\right)}=0$. Recalling from Eq. (16) that
$r_{b}^{(1)}=2M_{MS}^{\left(1\right)}\left(t,r_{b}^{(1)}\right),$ (37)
it follows from Eq. (35) that
$H_{1}^{2}\left(t\right)\left(r_{b}^{\left(1\right)}\right)^{3}-r_{b}^{\left(1\right)}+2M_{1}\left(t\right)=0.$ (38)
This equation is structurally analogous to the cubic Eq. (11). Its physically meaningful solutions, in agreement with (12a) and (12), are given by
$r_{b}^{\left(1\right)}\left(t\right)$ $=$ $\frac{2}{\sqrt{3}H_{1}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right),$
$r_{c}^{\left(1\right)}\left(t\right)$ $=$ $\frac{1}{H_{1}\left(t\right)}\cos\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right)$
$-\frac{1}{\sqrt{3}H_{1}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right)$
. (39b)
Therefore, within the adopted perturbative framework, we obtain analytical expressions for the temporal evolution, expressed in terms of a physical time coordinate, of both the black hole and cosmological apparent horizons as a result of the accretion process.
Moreover, the conditions required for the solutions in Eqs. (39) to be real are similar to those discussed in the previous section for Eqs. (12a) and (12), now for $M_{0}\rightarrow M_{1}\left(t\right)$ e $H_{0}\left(t\right)\rightarrow H_{1}\left(t\right)$.
Naturally, the analytical character of the solutions $r_{b}^{\left(1\right)}\left(t\right)$ and $r_{c}^{\left(1\right)}\left(t\right)$ depends on whether $M_{1}\left(t\right)$ and $H_{1}\left(t\right)$ can also be treated analytically. Therefore, we shall investigate the explicit solutions for specific $\rho_{ac}$ and $p_{ac}$.
IV Applications
IV.1 Dark energy accretion
Before analyzing the GCG, let us first consider a simpler case: the accretion of a fluid whose equation of state corresponds to that of dark energy, namely,
$p_{ac}=-\rho_{ac}\Rightarrow p_{\Lambda}=-\rho_{\Lambda}.$ (40)
From Eqs. (31), (33), and (36), we get
$F\left(t\right)$ $=$ $0,$ (41)
$H_{1}^{2}\left(t\right)$ $=$ $H_{0}^{2}+\frac{8\pi\rho_{\Lambda}}{3},$ (42)
$M_{1}\left(t\right)$ $=$ $M_{i}-\left(\frac{H_{0}^{2}}{2}+\frac{4\pi\rho_{\Lambda}}{3}\right)r_{0ac}^{3}.$ (43)
Therefore, from Eq. (35), we have
$M_{MS}^{\left(1\right)}\left(t,r_{b}^{\left(1\right)}\right)=M_{i}+\left(\frac{H_{0}^{2}}{2}+\frac{4\pi\rho_{\Lambda}}{3}\right)\left(\left(r_{b}^{\left(1\right)}\right)^{3}-r_{0ac}^{3}\right).$ (44)
A similar equation appears in Ref. [babichev2012backreaction], but in the case where $H_{0}=0$. In Ref. [babichev2012backreaction] treatment, Einstein’s equations are evaluated in a vacuum background, that is, $G_{\mu\gamma}\left[g_{\mu\gamma}^{(0)}\right]=0$. In contrast, our zeroth-order solution to Einstein’s equations is given by the McVittie equations.
IV.2 Generalized Chaplygin Gas
Let us now analyze the accretion of the GCG, whose equation of state is given by Eq. (6). From Eq. (7), we identify two distinct evolutionary phases of the universe. For small time scales, the universe is matter-dominated such that
$\rho^{(matter)}_{GCG}\approx\frac{B^{1/\left(1+\alpha\right)}}{a^{3}},$ (45)
while for large time scales, the universe behaves as a de Sitter spacetime, where
$\rho^{(dS)}_{GCG}\approx A^{1/\left(1+\alpha\right)}\Rightarrow p^{(dS)}_{GCG}=-\rho^{(dS)}_{GCG}$. (46)
Furthermore, there exists an effective equation of state in the intermediate regime between the matter-dominated and de Sitter phases, which can be obtained by expanding Eq. (7) as in Ref. [bento2002generalized]:
$\rho_{GCG}$ $\approx$ $A^{1/\left(1+\alpha\right)}+\left(\frac{1}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)}+$...,
$p_{GCG}$ $\approx$ $-A^{1/\left(1+\alpha\right)}+\left(\frac{\alpha}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)}+$...
Eqs. (47) are essential, since within the perturbative framework we are using, the contribution of the accreting fluid must be much smaller than those of the background fluid.
However, since we are assuming that the GCG drives the cosmic expansion, thus determining the background metric, while we are also interested in analyzing its accretion, we shall adopt the following strategy: we associate the dominant term in Eq. (47) with the background fluid characterized by $H_{0}$, and the remaining term with the perturbation.
For example, in a de Sitter universe:
$a\rightarrow\infty\Rightarrow\left\{\begin{array}[]{l}H_{0}=\sqrt{\frac{8\pi\rho^{(dS)}_{GCG}}{3}}=\sqrt{\frac{8\pi A^{1/\left(1+\alpha\right)}}{3}},\\ \rho_{ac}=\left(\frac{1}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)},\\ p_{ac}=\left(\frac{\alpha}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)}=\alpha\rho_{ac}.\end{array}\right$. (48)
For a matter-dominated universe:
$a^{(matter)}\Rightarrow\left\{\begin{array}[]{l}H_{0}=\sqrt{\frac{8\pi\rho^{(matter)}_{GCG}}{3}}=\frac{2}{3t},\\ \rho_{ac}=A^{1/\left(1+\alpha\right)},\\ p_{ac}=-A^{1/\left(1+\alpha\right)}=-\rho_{ac}.\end{array}\right.$. (49)
We shall now analyze each case separately.
IV.2.1 Matter dominated era
In a universe governed by the GCG during its matter-dominated phase, from Eq. (45) we have
$H_{0}\left(t\right)=\frac{2}{3t}.$ (50)
According to our perturbative construction, we shall assume that the accreting fluid obeys the subdominant term of the equation of state given in Eq. (47). Thus, its equation of state reads
$\rho_{ac}=-p_{ac}=A^{1/\left(1+\alpha\right)}.$ (51)
Consequently, from Eqs. (31), (33), and (36), we obtain
$F\left(t\right)$ $=$ $0,$ (52a)
$H_{1}^{2}\left(t\right)$ $=$ $\frac{4}{9t^{2}}+\frac{8\pi A^{1/\left(1+\alpha\right)}}{3},$ (52b)
$M_{1}\left(t\right)$ $=$ $M_{0}+M_{ac}+\left(H_{0}^{2}\left(t_{0ac}\right)-H_{1}^{2}\left(t\right)\right)\frac{r_{0ac}^{3}}{2},$
where, for an energy density given by (51), $M_{ac}=4\pi r_{ac}^{3}\rho_{ac}/3$. Therefore, the effective mass expressed as a first-order perturbation as given by Eq. (35), becomes
$M_{MS}^{\left(1\right)}=M_{i}+\left(\frac{2}{9t^{2}}+\frac{4\pi A^{1/(1+\alpha)}}{3}\right)\left(\left(r_{b}^{\left(1\right)}\right)^{3}-r_{0ac}^{3}\right).$ (53)
We now study how accretion affects the evolution of the cosmological and black hole apparent horizons in comparison with the non-accreting case. To this end, we compare Eqs. (12a) and (12) with Eqs. (39).
From Eqs. (12a) and (12), as discussed in section II, when the equation
$3\sqrt{3}M_{0}H_{0}\left(t\right)=1$ (54)
is satisfied, this corresponds to the instant at which the cosmological and black hole horizons form, being equal at that moment. For
$3\sqrt{3}M_{0}H_{0}\left(t\right)<1,$ (55)
the horizons become distinct.
Hence, in a matter-dominated universe where the Hubble parameter is given by Eq. (50), the horizons form at the instant
$t_{0}=2\sqrt{3}M_{0}$ (56)
and become distinct for $t>t_{0}$.
Analogously, we can identify the instant $t_{0ac}$ corresponding to the formation of the horizons due to accretion. Similarly to the non-accreting case, from Eqs. (39) it follows that the horizons are formed at instant $t_{0ac}$ satisfying
$3\sqrt{3}M_{1}\left(t_{0ac}\right)H_{1}\left(t_{0ac}\right)=1.$ (57)
Using Eqs. (52b) and (LABEL:M1_matter), Eq. (57) admits, as its positive solution,
$t_{0ac}=\frac{t_{0}}{\sqrt{1-72\pi M_{0}^{2}A^{1/\left(1+\alpha\right)}}}=\frac{t_{0}}{\sqrt{1-72\pi M_{0}^{2}\rho_{ac}}},$ (58)
which, in turn, imposes the bound $A<\left(72\pi M_{0}^{2}\right)^{-\left(1+\alpha\right)}$.
It is straightforward to notice that, for $A=0$, that is, when $\rho_{ac}=0$, $t_{0ac}=t_{0}$. However, it is important to emphasize that the delay in the formation time of the apparent horizons is not due to the accretion process itself, but rather to the fact that there is now more matter available for accretion. More precisely, this refers to matter with mass $M_{ac}=4\pi r_{0ac}^{3}\rho_{ac}/3$.
From this, we can clearly observe the effect of available matter for accretion on the instant of horizon formation by plotting Eq. (58), as shown in Fig. 1. The plot shows that, as $A$ increases, the black hole and cosmological horizons takes longer to form.
Figure 1: Apparent black hole and cosmological horizon formation time as a function of AA (solid lines), as given by (58). The dashed line represent the instant of formation of the apparent horizons without available matter for accretion, $t_{0}$, given by 56. The mass parameter $M_{0}$ is set to 11. We also consider $\alpha=0.1$.
This result appears counterintuitive at first. For instance, in static black holes such as the Schwarzschild solution, a greater amount of matter typically leads to faster gravitational collapse and earlier horizon formation. To clarify this result, we now examine the evolution of the horizons.
The horizons obtained from Eqs. (39), together with Eqs. (52b) and (LABEL:M1_matter), are shown in Fig. 2 for different values of $A$. As illustrated, larger values of $A$, and therefore higher energy densities of matter available to the accretion process, cause the black hole horizon to expand, starting from an initial radius of $3M_{0}$ and settling at constant values greater than the Schwarzschild radius $2M_{0}$ after a sufficiently long time. In contrast, the cosmological horizon decreases with increasing $A$, eventually reaching a steady monotonic regime.
Figure 2: Temporal evolution of the apparent black hole horizon without available matter for accretion (solid black line) and with available matter for accretion (colored solid lines); and of the cosmological horizons without available matter for accretion (black dashed line) and with available matter for accretion (colored dashed lines) for different values of $A$. The mass parameter $M_{0}$ is set to $M_{0}=1$, such that $t$ is measured in units of $M_{0}$. Also, we consider $\alpha=0.1$.The horizontal dashed line represents the asymptote $r=2M_{0}$. The vertical solid line represents the instant $t_{0}$ at which the horizons form in the case without accretion, as given by Eq. (56).
Let us now examine the horizons in their asymptotic stage, beginning with the case without accretion. After a sufficiently long time, such that $t>>t_{0ac}$, the horizons given by Eqs. (12a) and (12) behave as
$r_{b}^{\left(0\right)}\left(t\right)\approx 2M_{0}\text{, \ }r_{c}^{\left(0\right)}\left(t\right)\approx\frac{1}{H_{0}}-M_{0}\approx\frac{3t}{2}-M_{0}.$ (59)
Similarly, for Eqs. (39), we find that for time scales such that $t>>t_{0ac}$, the behavior of the horizons in the presence of accretion is given by
$r_{b}^{\left(1\right)}\left(t\right)$ $\approx$ $2M_{1}\left(t\right)\approx 3M_{0}-36\pi A^{1/(1+\alpha)}M_{0}^{3},$ (60a)
$r_{c}^{\left(1\right)}\left(t\right)$ $\approx$ $\frac{1}{H_{1}\left(t\right)}-M_{1}\left(t\right)\approx\left(\frac{4}{9t^{2}}+\frac{4\pi A^{1/(1+\alpha)}}{3}\right)^{-1/2}$ (60b)
$-\frac{3M_{0}}{2}+18\pi A^{1/(1+\alpha)}M_{0}^{3},$
where we used the fact that $r_{0ac}\equiv r_{b}^{(1)}(t_{0ac})=3M_{0}$.
For the black hole horizon with accretion to be larger than in the case without accretion, the condition $A<\left(12\pi M_{0}^{2}\right)^{-\left(1+\alpha\right)}$ must be satisfied, according to Eq. (60a). However, due to the constraint previously imposed by Eq. (58), this requirement is automatically fulfilled.
Fig. 2 also shows that, for values of $A$ on the order of $10^{-3}{$\mathrm{m}$}^{-2(1+\alpha)}$, the evolution of the horizons differs significantly from the case without accretion. Therefore, we assume that our perturbative scheme remains valid for an energy density of the accreting fluid of the order of $10^{-4}{$\mathrm{m}$}^{-2(1+\alpha)}$ when compared to that of the background fluid. Therefore, given Eq. (49), this implies that our perturbative approach remains valid only for timescales shorter than
$t=10^{2}\sqrt{\frac{32\pi}{27\rho_{ac}}}.$ (61)
Furthermore, it is worth noting that the values of the Chaplygin parameter $A$ explored here are chosen for illustrative clarity and exceed those expected from cosmological considerations. Observationally consistent models typically give $10^{-60}{$\mathrm{m}$}^{-2(1+\alpha)}<A<10^{-54}{$\mathrm{m}$}^{-2(1+\alpha)}$ for $\alpha\leq 0.1$, as in Refs. [aurich2018concordance, barreiro2008wmap]. Using enhanced values in our simulations makes the accretion effects numerically visible while preserving the qualitative behavior of the horizon evolution. The physically motivated range of $A$ would simply place these effects at correspondingly smaller amplitudes, without modifying the trends highlighted by our results.
Therefore, although the apparent horizon of the black hole attains larger values due to accretion while the cosmological apparent horizon decreases, as expected, we obtained the counterintuitive result given by Eq. (58), namely that the greater the amount of accreting matter, the shorter the time required for the black hole horizon to form.
For a cosmologically coupled black hole like the McVittie solution, the backreaction of accretion on the metric alters this behavior. The inflow of accreting matter locally increases the energy density and consequently reduces the rate at which cosmic expansion decelerates, as shown in Eqs. (60a) and (52b). The resulting higher effective expansion rate counteracts the growth of the local gravitational term, namely, the apparent black hole horizon, as expressed in Eqs. (60a) and (LABEL:M1_matter), thereby delaying the instant at which the apparent horizon condition is satisfied.
IV.2.2 de Sitter era
In a de Sitter–type universe governed by the GCG, the Hubble parameter is given by
$H_{0}=\sqrt{\frac{8\pi A^{1/(1+\alpha)}}{3}}\equiv H_{A}.$ (62)
The cosmological scale factor, in turn, takes the form
$a\left(t\right)=ke^{H_{A}t},$ (63)
where $k$ is a positive constant.
According to our perturbative approach, since the background is determined by an equation of state of the form $\rho^{(dS)}_{GCG}=-p^{(dS)}_{GCG}=A^{1/\left(1+\alpha\right)}$, we assume that the accreting fluid corresponds to the subdominant term in Eq. (47), namely
$p_{ac}=\alpha\rho_{ac}=\left(\frac{\alpha}{1+\alpha}\right)\frac{B}{A^{\alpha/(1+\alpha)}}a^{-3\left(1+\alpha\right)}.$ (64)
In this case, the flux $F\left(t\right)$ defined in Eq. (31) is not, a priori, equal to zero. Hence, the radial velocity $v$ must be obtained from the continuity equation
$\left(u^{\mu}\right)^{(0)}\partial_{\mu}\rho_{ac}+\left(\rho_{ac}+p_{ac}\right)\nabla_{\mu}\left(u^{\mu}\right)^{(0)}=0.$ (65)
It is important to note that, within the adopted approximation, the continuity equation is evaluated in the background metric, i.e. $u^{\mu}\rightarrow\left(u^{\mu}\right)^{\left(0\right)}$. Furthermore, both $\rho_{ac}$ and $p_{ac}$ depend only on time, and the accretion is evaluated at the black hole horizon, such that $r=r_{b}^{\left(0\right)}\left(t\right)$ and $g_{00}^{\left(0\right)}=0$. Additionally, in a dark energy-dominated era, the Hubble parameter is constant and, consequently, the horizon radius is also constant, given by $r_{b}^{\left(0\right)}=2M_{0}$.
A full analytical solution of the continuity equation in this setting is possible in principle but offers limited physical insight. Instead, we focus on the regime in which the radial velocity satisfies $v>>M_{0}H_{A}$, a limit that is both physically relevant for the systems under consideration and leads to expressions that more transparently capture the underlying dynamics.
As discussed previously, Refs. [aurich2018concordance, barreiro2008wmap] typically give $10^{-60}{$\mathrm{m}$}^{-2(1+\alpha)}<A<10^{-54}{$\mathrm{m}$}^{-2(1+\alpha)}$ for $\alpha\leq 0.1$. If we conservatively assume an initial accretion velocity $v_{0}=0.01c_{s}$, where $c_{s}=\sqrt{{\partial p}/{\partial\rho}}$ is the sound speed derived from the equation of state (64), we obtain $v_{0}=0.01\alpha^{1/2}$.
Hence, the condition $v>>M_{0}H_{A}$ is satisfied for black holes whose mass $M_{0}$ fulfills $M_{0}<<v_{0}/H_{A}$. For $\alpha=0.1$ and $A=10^{-54}{$\mathrm{m}$}^{-2(1+\alpha)}$, this yields $M_{0}<<3.85\times 10^{21}{$\mathrm{m}$}$. In the units adopted here, a supermassive black hole with a mass of one billion solar masses corresponds to $M_{0}\sim 10^{12}{$\mathrm{m}$}$, indicating that the inequality $M_{0}<<v_{0}/H_{A}$ is easily satisfied.
However, since the full time dependence of $v$ is not known, we assume that this approximation holds only over astrophysical timescales. According to Eq. (64), the density of the accreting fluid decreases sharply with cosmological time, indicating that the accretion dynamics at cosmological timescales would be significantly altered.
Therefore, from Eq. (31) we get
$F\left(t\right)$ $=$ $64\pi H_{A}^{2}M_{0}^{4}(1+\alpha)\int_{t_{0ac}}^{t}\rho_{ac}d\bar{t}=$ (66)
$=$ $\frac{64\pi H_{A}M_{0}^{4}}{3}\left(\rho_{ac0}-\rho_{ac}\left(t\right)\right)$
where $\rho_{ac0}$ denotes the density given by Eq. (64) at the instant $t_{0ac}$ when accretion begins.
For the function $H_{1}\left(t\right)$ defined in Eq. (33), we have
$H_{1}\left(t\right)=\left(\frac{8\pi}{3}\right)^{1/2}\sqrt{A^{1/(1+\alpha)}-\alpha\rho_{ac}}.$ (67)
The expression above imposes the constraint
$A^{1/(1+\alpha)}>\alpha\rho_{ac},$ (68)
as expected within our perturbative framework, since the left-hand side is associated with the background density, whereas the right-hand side corresponds to the accreting fluid density.
Solving inequality (68) for $t$, we obtain
$t>\ln\left[\left(\frac{\alpha}{1+\alpha}\frac{B}{A}k^{-3/\left(1+\alpha\right)}\right)^{\frac{1+\alpha}{3H_{A}}}\right].$ (69)
Thus, we set the instant when accretion begins in this phase of the universe as
$t_{0ac}=\ln\left[\left(\frac{\alpha}{1+\alpha}\frac{B}{A}k^{-3/\left(1+\alpha\right)}\right)^{\frac{1+\alpha}{3H_{A}}}\right].$ (70)
This equation allows us to relate the horizon formation time to different values of $A$. As shown in Fig. 3, plotting $t_{0ac}$ reveals that, for smaller values of $A$, the horizons take longer to form.
This result is the opposite of that obtained from Eq. (58). We can interpret this behavior by recalling that, during the matter-dominated era, the parameter $A$ is directly associated with the subdominant term, namely, the energy density of the accreting fluid. In contrast, in a de Sitter universe, $A$ is directly related to the background energy density. Consequently, a smaller background energy density implies less fluid available for accretion, which in turn delays the formation of the apparent horizons.
Figure 3: Apparent black hole and cosmological horizon formation time as a function of $A$, as given by (70). We consider $\alpha=0.1$, B=1AB=1-A and $k=0.39$.
Finally, for the function $M_{1}\left(t\right)$ defined in Eq. (36), we obtain
$M_{1}\left(t\right)$ $=$ $M_{0}+\frac{32\pi M^{3}_{0}}{3}\rho_{ac0}+18\pi\alpha M_{0}^{3}\rho_{ac}$ (71)
$+\frac{64H_{A}M_{0}^{4}\pi}{3}\left(\rho_{ac0}-\rho_{ac}\left(t\right)\right),$
where we used $M_{ac}(t_{0ac})=4\pi r^{3}_{0ac}\rho_{ac0}/3$.
Therefore, since at large times $r_{b}^{\left(1\right)}\approx 2M_{1}\left(t\right)$ and $r_{c}^{\left(1\right)}\approx\left(1/H_{1}\left(t\right)\right)-M_{1}\left(t\right)$, and given that $\rho_{ac}\left(t\right)$ decreases with time according to Eq. (64), the black hole horizon becomes larger in the presence of accretion than in its absence, while the cosmological horizon becomes correspondingly smaller.
V Conclusions
In this work, we investigated the accretion of cosmic dark fluids, particularly the GCG, onto cosmologically coupled black holes described by the McVittie metric. By employing a perturbative formalism that takes into account the backreaction of accretion on the metric components, we derived analytical expressions for the effective black hole mass and for the evolution of both the black hole and cosmological apparent horizons. This framework allowed us to analyze how the local accretion process interacts with the global cosmological background in different epochs of the universe.
In the matter-dominated era, our results reveal a somewhat counterintuitive behavior. One might expect that the presence of additional matter available for accretion would accelerate the black hole formation process. However, due to the backreaction effects inherent to the McVittie metric, the situation is reversed. The inflow of matter locally increases the energy density, which in turn reduces the deceleration of the cosmic expansion. As a result, the effective Hubble parameter becomes larger, counteracting the local gravitational collapse. This dynamic causes the apparent horizon of the black hole to form later for higher values of $A$, that is, when the available matter for accretion is greater.
The analysis of Eq. (58) confirms that the horizon formation time $t_{0ac}$ increases with $A$, which means that the cosmological coupling and the backreaction jointly act to delay the black hole horizon formation compared to the non-accreting case.
In contrast, during the de Sitter phase, the behavior of horizon formation exhibits the opposite dependence on $A$. In this regime, the parameter $A$ is directly tied to the background energy density of the universe rather than to the accreting fluid. Consequently, a smaller value of $A$ corresponds to a lower cosmological energy density, implying less available fluid to be accreted. This reduced inflow slows down the local mass growth, thereby delaying the formation of the horizons. Hence, while in the matter-dominated universe a higher $A$ delays the formation time due to backreaction, in the de Sitter regime, it is a smaller $A$ that produces a similar delay. This complementary behavior highlights the subtle interplay between local and global dynamics in cosmologically coupled systems.
Overall, since the McVittie metric provides a bridge between cosmological expansion and local gravitational collapse, our formalism could be particularly relevant for the study of primordial black holes, whose formation and evolution are naturally embedded in an expanding universe. Understanding how accretion and backreaction operate in such early-universe scenarios could shed light on the role of cosmologically coupled black holes in structure formation and in the broader dynamics of the dark sector.
Acknowledgements
The work of O.B. is partially supported by FCT (Fundação para a Ciência e Tecnologia, Portugal) through the project 2024.00252.CERN. The work of L.F.R. is supported by FCT through the grant 2024.06413.BD. Both authors are partially supported by UID/04650 - Centro de Física das Universidades do Minho e do Porto. M. C. B. thanks the University of Porto for hosting part of this study.

### v2

@@
 For the matter-dominated era, the analytical expression shows that the greater the amount of matter available for accretion, the longer the accretion takes to start.
-I Introduction
+1 Introduction
 […]
 While the cosmological constant remains the simplest and most widely adopted model for dark energy, a variety of dynamical models have been proposed to address the theoretical and observational challenges.
-Among these, the Generalized Chaplygin Gas (GCG) has attracted attention due to its ability to interpolate between dust-like matter at early times and a dark energy–dominated regime at late times [kamenshchik2001alternative, Bilic:2001cg, bento2002generalized].
-The GCG model not only provides a compelling observational framework to account for the universe’s accelerated expansion and to address the dark matter problem [bento2003wmap, aurich2018concordance, barreiro2008wmap], but also offers promising avenues for addressing fundamental issues such as the cosmological constant problem [bertolami2023seeding], the Hubble tension [yang2019dawn], the Swampland conjectures [bertolami2024sitter], and inflation through a model inspired on the GCG [Bertolami:2006zg].
+Among these, the Generalized Chaplygin Gas (GCG) has attracted attention due to its ability to interpolate between dust-like matter at early times and a dark energy-dominated regime at late times [8, 9, 10].
+The GCG model not only provides a compelling observational framework to account for the universe’s accelerated expansion and to address the dark matter problem [11, 12, 13], but also offers promising avenues for addressing fundamental issues such as the cosmological constant problem [14], the Hubble tension [15], the Swampland conjectures [16], and inflation through a model inspired by the GCG [17].
 Despite significant progress in constraining dark energy models, its interaction with compact astrophysical objects remains a relatively unexplored frontier.
@@
 Some studies have already explored dark energy and the GCG in the context of compact objects, such as in Refs.
-[Mazur:2001fv, bertolami2005chaplygin, gorini2009more].
+[18, 19, 20].
 Other works have also investigated dark energy and the GCG in accretion scenarios, such as in Refs.
-[babichev2005accretion, rodrigues2012accretion], although there remains substantial room for further development in this area.
-Classical accretion theory has been instrumental in describing the infall of baryonic matter and the associated electromagnetic signatures [bondi1952spherically, petterson1980variations, michel1972accretion, porth2017black].
+[21, 22], although there remains substantial room for further development in this area.
+Classical accretion theory has been instrumental in describing the infall of baryonic matter and the associated electromagnetic signatures [23, 24, 25, 26].
 However, in the context of a universe dominated by a dark energy, it becomes imperative to consider whether and how black holes may accrete this non-standard component.
 The accretion of dark energy, particularly when modeled as a perfect fluid with negative pressure, introduces novel theoretical possibilities.
-In scenarios involving phantom energy, that satisfies the condition $p+\rho<-1$, it has been shown that the accretion process could lead to a decrease in black hole mass, a counterintuitive result that challenges conventional views of black hole growth and thermodynamics [babichev2004black].
+In scenarios involving phantom energy, that satisfies the condition $p+\rho<-1$, it has been shown that the accretion process could lead to a decrease in black hole mass, a counterintuitive result that challenges conventional views of black hole growth and thermodynamics [27].
 However, despite the consistency of such approaches, certain aspects are often neglected, largely due to the considerable complexity involved in developing a comprehensive accretion model.
@@
 In the standard test-fluid approximation, where the accreting material is assumed too small to significantly modify the black hole’s spacetime and the geometry is treated as fixed, the accreting matter or energy is treated as having negligible influence on the background spacetime geometry.
+Backreaction effects have been explored in steady accretion models where the self-gravity of the accreting fluid modifies the background geometry; for example, the fully relativistic treatment in [28] examines how the gravitational influence of matter affects accretion compared to the Bondi limit, and Ref.
+[29] studies steady critical accretion effects on the characteristics of the sonic points.
 However, for accretion processes that persist over cosmological timescales or involve significant energy fluxes, the assumption of a fixed black hole metric may no longer be valid.
@@
 In other words, we seek to study a black hole whose spacetime is not asymptotically flat, but rather asymptotically approaches a flat Friedmann-Lemaitre-Robertson-Walker (FLRW) cosmology.
-For this purpose, we adopt the McVittie metric [mcvittie1933mass].
+For this purpose, we adopt the McVittie metric [30].
 This spacetime solution has been the subject of ongoing debate for nearly eighty years regarding its physical consistency in describing black holes embedded in an expanding universe.
-However, recent studies have demonstrated that the McVittie metric not only provides a consistent framework for modeling black holes that are asymptotically FLRW [kaloper2010mcvittie, lake2011more, nandra2012effect, nandra2012effect1, faraoni2015cosmological], but also serves as an important foundational model for the development of more sophisticated and realistic metrics to this end [gao2011black].
+However, recent studies have demonstrated that the McVittie metric not only provides a consistent framework for modeling black holes that are asymptotically FLRW [31, 32, 33, 34, 35], but also serves as an important foundational model for the development of more sophisticated and realistic metrics to this end [36].
 The present work does not aim to delve into the broader debate concerning the validity and limitations of the McVittie metric.
 Nevertheless, since this metric will be employed in the next section to model both the accretion process and the associated backreaction, the results obtained may offer valuable insights and contribute to future efforts toward constructing more complete metrics capable of describing astrophysical objects within an expanding universe.
-This paper is organized as follows: In Section II, We analyze the McVittie metric, with an emphasis on the effective mass associated with this geometry and on the time-dependent solutions of the apparent horizons.
-In Section III, we use a perturbative analysis that has already been developed that accounts for backreaction effects to study accretion within the McVittie spacetime.
-In Section IV, we apply the results developed in Section III to the GCG, focusing on two distinct analyses: one corresponding to a matter-dominated universe and the other to a de Sitter universe.
-Finally, in Section V, we present our conclusions and discuss the implications of the results.
+This paper is organized as follows: In Section 2 we analyze the McVittie metric, with an emphasis on the effective mass associated with this geometry and on the time-dependent solutions of the apparent horizons.
+In Section 3, we use a perturbative analysis that has already been developed that accounts for backreaction effects to study accretion within the McVittie spacetime.
+In Section 4, we apply the results developed in Section 3 to the GCG, focusing on two distinct analyses: one corresponding to a matter-dominated universe and the other to a de Sitter universe.
+Finally, in Section 5, we present our conclusions and discuss the implications of the results.
 Throughout this work we use units in which $c=G=1$ and adopt the metric signature to be +++-+++. Also, we will use the notation $^{\cdot}=\partial/\partial t$ and $^{\prime}=\partial/\partial r$.
-II The McVittie metric and the GCG
-The McVittie metric can be expressed in several equivalent forms; however, in this work, we adopt the formulation presented in Ref. [kaloper2010mcvittie]. The line element is given by
+2 The McVittie metric and the GCG
+The McVittie metric can be expressed in several equivalent forms; however, in this work, we adopt the formulation presented in Ref. [31]. The line element is given by
 $ds^{2}=-\left(1-\frac{2M_{0}}{r}-H_{0}^{2}\left(t\right)r^{2}\right)dt^{2}-\frac{2rH_{0}\left(t\right)}{\sqrt{1-\frac{2M_{0}}{r}}}dtdr+\frac{dr^{2}}{1-\frac{2M_{0}}{r}}+r^{2}d\Omega_{2}^{2},$ (1)
@@
 The Einstein field equations for metric (1) yield the following relations:
-$-3H_{0}^{2}(t)$ $=$ $8\pi\left(T_{0}^{\text{ }0}\right)^{\left(0\right)},$ (2a)
-$-3H_{0}^{2}-\frac{2\dot{H}_{0}}{\sqrt{1-\frac{2M_{0}}{r}}}$ $=$ $8\pi\left(T_{1}^{\text{ }1}\right)^{\left(0\right)}.$ (2b)
+$8\pi\left(T_{0}^{\text{ }0}\right)^{\left(0\right)}$ $=-3H_{0}^{2}(t),$ (2)
+$8\pi\left(T_{1}^{\text{ }1}\right)^{\left(0\right)}$ $=-3H_{0}^{2}-\frac{2\dot{H}_{0}}{\sqrt{1-\frac{2M_{0}}{r}}}.$ (3)
 If we define $\left(T_{\nu}^{\text{ }\mu}\right)^{\left(0\right)}$ such that $\left(T_{0}^{\text{ }0}\right)^{\left(0\right)}=-\rho^{\left(0\right)}$ and $\left(T_{1}^{\text{ }1}\right)^{\left(0\right)}=p^{\left(0\right)}$, where $\rho^{\left(0\right)}$ and $p^{\left(0\right)}$ are the energy density and isotropic pressure, respectively, Eq.
-(2a) can be readily recognized as the Friedmann equation.
+(2) can be readily recognized as the Friedmann equation.
 This indicates that the presence of a black hole does not affect the homogeneity of the energy density distribution.
 However, Eq.
-(2b) reveals that the same is not true for the pressure, which acquires a spatial dependence due to the black hole’s influence.
+(3) reveals that the same is not true for the pressure, which acquires a spatial dependence due to the black hole’s influence.
 To further elucidate the non-homogeneous nature of the pressure, we consider the energy conservation equation derived from the Einstein field equations. This relation is given by
@@
 The authors in Refs.
-[nandra2012effect, nandra2012effect1] have shown that for a matter-dominated universe, it is appropriate to take $f\left(t\right)=\rho^{\left(0\right)}$, such that the pressure can be written as
+[33, 34] have shown that for a matter-dominated universe, it is appropriate to take $f\left(t\right)=\rho^{\left(0\right)}$, such that the pressure can be written as
 $p^{\left(0\right)}=\frac{\rho^{\left(0\right)}}{\sqrt{1-\frac{2M_{0}}{r}}}-\rho^{\left(0\right)}.$ (6)
@@
 where $A>0$ and $0<\alpha\leq 1$.
-The Chaplygin gas corresponds to the choice $\alpha=1$ [kamenshchik2001alternative, bento2002generalized, bento2003wmap].
+The Chaplygin gas corresponds to the choice $\alpha=1$ [8, 10, 11].
 In the framework of FLRW cosmology, the relativistic energy conservation associated to the equation of state (7) leads to an evolving energy density:
@@
 Other key aspects of the McVittie metric relevant to our analysis are the effective mass and the apparent horizon.
-The effective mass of the black hole can be computed using the Misner-Sharp mass function $M_{MS}^{\left(0\right)}$ [misner1964relativistic], defined by
+The effective mass of the black hole can be computed using the Misner-Sharp mass function $M_{MS}^{\left(0\right)}$ [37], defined by
 $1-\frac{2M_{MS}^{\left(0\right)}}{r}=\left(g^{\mu\nu}\right)^{\left(0\right)}\partial_{\mu}r\partial_{\nu}r,$ (10)
-which, using Eqs. (1) and (2a), results in
+which, using Eqs. (1) and (2), results in
 $M_{MS}^{\left(0\right)}\left(t,r\right)=M_{0}+\frac{H_{0}^{2}r^{3}}{2}.$ (11)
-The apparent horizon of the black hole, denoted $r_{h}^{\left(0\right)}$, can be determined from the condition $g_{00}^{\left(0\right)}=0$, which, using Eq. (1), leads to the following cubic equation:
-$H_{0}^{2}\left(r_{h}^{\left(0\right)}\right)^{3}-r_{h}^{\left(0\right)}+2M_{0}=0.$ (11)
-As shown in Refs. [nickalls1993new, faraoni2012making] the solutions of the cubic equation above can be expressed as:
-$r_{1}\left(t\right)$ $=$ $\frac{2}{\sqrt{3}H_{0}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{0}H_{0}\left(t\right)\right)}{3}\right),$ (12a)
+Since the McVittie spacetime is dynamical according to Eq.
+(11), a global event horizon cannot, in general, be identified without knowledge of the full future development of the spacetime.
+It is therefore more appropriate to characterize black hole boundaries quasi-locally in terms of apparent horizons.
+In a given spacelike hypersurface, an apparent horizon is defined as the outermost marginally outer trapped surface (MOTS).
+One may equivalently characterize the apparent horizon in terms of the optical scalars—namely, the expansion $\Theta$, the twist $\omega$, and the shear $\sigma$.
+In this framework, the apparent horizon is specified by the conditions $\Theta_{+}=0$, $\Theta_{-}<0$, $\omega=0$, and $\sigma=0$, where $\Theta_{+}=0$ corresponds to the vanishing expansion of the outgoing null congruence, while $\Theta_{-}<0$ ensures that the ingoing null congruence remains converging [28].
+For a spherically symmetric spacetime given by the metric (1), the symmetry two-spheres have area $\mathcal{S}=4\pi r^{2}$ and, in the coordinates used here, the outgoing and ingoing expansions take the form
+$\Theta_{\pm}=\frac{2}{r}\frac{dr}{ds_{\pm}}\propto\sqrt{1-\frac{2M_{0}}{r}}\left(H_{0}r\pm\sqrt{1-\frac{2M_{0}}{r}}\right).$ (12)
+Therefore, the marginally trapped condition $\Theta_{+}=0$ yields
+$H_{0}^{2}\left(r_{h}^{\left(0\right)}\right)^{3}-r_{h}^{\left(0\right)}+2M_{0}=0,$ (13)
+where $r_{h}^{\left(0\right)}$ represents the geometrical locus of the apparent horizon.
+As shown in Refs. [38, 39] the solutions of the cubic equation above can be expressed as:
+$r_{1}\left(t\right)$ $=\frac{2}{\sqrt{3}H_{0}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{0}H_{0}\left(t\right)\right)}{3}\right),$ (14)
 $r_{2}\left(t\right)$ $=\frac{1}{H_{0}\left(t\right)}\cos\left(\frac{\arcsin\left(3\sqrt{3}M_{0}H_{0}\left(t\right)\right)}{3}\right)$
@@
 As discussed in Ref.
-[faraoni2012making], the solution $r_{3}$ can be discarded as it is unphysical, since it is always negative.
+[39], the solution $r_{3}$ can be discarded as it is unphysical, since it is always negative.
 The solution $r_{1}$ can be interpreted as the apparent horizon of the black hole, as it reduces to the Schwarzschild radius $r_{1}\rightarrow 2M_{0}$ in the limit $H_{0}\rightarrow 0$.
@@
 Furthermore, the solutions given by Eqs.
-(12a) and (12) impose the following conditions: for $3\sqrt{3}M_{0}H_{0}<1$, the horiozons $r_{b}$ and $r_{c}$ are distinct; when $3\sqrt{3}M_{0}H_{0}=1$, the black hole horizon coincides with the cosmological horizon; and when $3\sqrt{3}M_{0}H_{0}>1$, no horizons exist and the black hole becomes a naked singularity.
+(14)-(2) impose the following conditions: for $3\sqrt{3}M_{0}H_{0}<1$, the horizons $r_{b}$ and $r_{c}$ are distinct; when $3\sqrt{3}M_{0}H_{0}=1$, the black hole horizon coincides with the cosmological horizon; and when $3\sqrt{3}M_{0}H_{0}>1$, no horizons exist and the black hole becomes a naked singularity.
 Finally, it is also important to compute the expansion velocity of the apparent horizons. Differentiating (13) with respect to time yields the evolution velocity of the apparent horizon:
@@
 In the following section, we shall build upon the results presented thus far to construct a more refined model that accounts not only for accretion but also for the backreaction effects on the spacetime geometry.
-III Perturbatitve approach to backreaction
-The process of accretion within the framework of the McVittie metric has been investigated in previous works [gao2011black, gao2008does, guariento2012realistic].
+3 Perturbative approach to backreaction
+The process of accretion within the framework of the McVittie metric has been investigated in previous works [36, 40, 41].
 The approach adopted in these studies primarily consists in promoting the mass parameter $M_{0}$ in Eq.
@@
 To investigate this backreaction, we employ a perturbative method as proposed in Ref.
-[babichev2012backreaction].
-We begin by assuming that the backreaction induces both a temporal and radial dependence of the mass, as well as an inhomogeneity in the Hubble parameter. Consequently, the metric (1) becomes:
-$ds^{2}=-\left(1-\frac{2M\left(t,r\right)}{r}-H\left(t,r\right)^{2}r^{2}\right)dt^{2}-\frac{2rH\left(t,r\right)}{\sqrt{1-\frac{2M\left(t,r\right)}{r}}}dtdr+\frac{dr^{2}}{1-\frac{2M\left(t,r\right)}{r}}+r^{2}d\Omega_{2}^{2}.$ (14)
+[42].
+We do it so by assuming that accretion induces both a temporal and radial dependence of the mass, as well as an inhomogeneity in the Hubble parameter.
+Consequently, the metric (1) becomes:
+$ds^{2}$ $=$ $-\left(1-\frac{2M\left(t,r\right)}{r}-H\left(t,r\right)^{2}r^{2}\right)dt^{2}-\frac{2rH\left(t,r\right)}{\sqrt{1-\frac{2M\left(t,r\right)}{r}}}dtdr$ (18)
+$+$ $\frac{dr^{2}}{1-\frac{2M\left(t,r\right)}{r}}+r^{2}d\Omega_{2}^{2}.$
 The Einstein equations for the line element (18) then take the following form:
-$8\pi\mathcal{T}_{0}^{\text{ }0}$ $=$ $-3H^{2}-2rHH^{\prime}-\frac{2M^{\prime}}{r^{2}},$ (15a)
-$8\pi\mathcal{T}_{0}^{\text{ }1}$ $=$ $2rH\dot{H}+\frac{2\dot{M}}{r^{2}},$ (15b)
-$8\pi\mathcal{T}_{1}^{\text{ }0}$ $=$ $-\frac{2\dot{M}}{\left(r-2M\right)^{2}},$ (15c)
-$8\pi\mathcal{T}_{1}^{\text{ }1}$ $=$ $-\frac{1}{r^{2}\left(r-2M\right)^{2}}\left[3r^{2}H^{2}\left(r-2M\right)^{2}\right$.
+$8\pi\mathcal{T}_{0}^{\text{ }0}$ $=-3H^{2}-2rHH^{\prime}-\frac{2M^{\prime}}{r^{2}},$ (19)
+$8\pi\mathcal{T}_{0}^{\text{ }1}$ $=2rH\dot{H}+\frac{2\dot{M}}{r^{2}},$ (20)
+$8\pi\mathcal{T}_{1}^{\text{ }0}$ $=-\frac{2\dot{M}}{\left(r-2M\right)^{2}},$ (21)
+$8\pi\mathcal{T}_{1}^{\text{ }1}$ $=-\frac{1}{r^{2}\left(r-2M\right)^{2}}\left[3r^{2}H^{2}\left(r-2M\right)^{2}\right$. (22)
 $\left.+2\left(r-2M\right)\left(\left(r-2M\right)M^{\prime}+r^{3}\sqrt{1-\frac{2M}{r}}\dot{H}\right)\right$.
 $\left.+2r^{3}H\left(\left(r-2M\right)^{2}H^{\prime}+\sqrt{1-\frac{2M}{r}}\dot{M}\right)\right].$
-One also finds that the Einstein tensors $G_{2}^{\text{ }2}=G_{3}^{\text{ }3}\neq 0$; however, these components yield no independent equations, as they follow from Eqs. (15a)-(15c).
+One also finds that the Einstein tensors $G_{2}^{\text{ }2}=G_{3}^{\text{ }3}\neq 0$; however, these components yield no independent equations, as they follow from Eqs. (19)-(21).
 If we define the function
 $\mathcal{M}_{MS}\left(t,r\right)=\frac{r}{2}\left(1-g_{00}\right)=M\left(t,r\right)+\frac{H\left(t,r\right)^{2}r^{3}}{2},$ (23)
-where $g_{\mu\nu}$ is the metric associated with the line element (14), then Eqs. (15a) and (15b) can be rewritten as
+where $g_{\mu\nu}$ is the metric associated with the line element (18), then Eqs. (19) and (20) can be rewritten as
 $\mathcal{M}_{MS}^{\prime}\left(t,r\right)=\mathcal{B}\left(t,r\right)$ (24)
 and
-[…]
+$\dot{\mathcal{M}}_{MS}\left(t,r\right)=\mathcal{A}\left(t,r\right)$ (25)
 where
@@
 The system of partial differential equations (24)-(25) admits the general solution
-$\mathcal{M}_{MS}\left(t,r\right)$ $=$ $M_{i}+\int_{t_{0ac}}^{t}\mathcal{A}\left(\bar{t},r_{0ac}\right)~d\bar{t}+\int_{r_{0ac}}^{r_{b}}\mathcal{B}\left(t,\bar{r}\right)~d\bar{r},$
+$\mathcal{M}_{MS}\left(t,r\right)=M_{i}+\int_{t_{0ac}}^{t}\mathcal{A}\left(\bar{t},r_{0ac}\right)~d\bar{t}+\int_{r_{0ac}}^{r_{b}}\mathcal{B}\left(t,\bar{r}\right)~d\bar{r},$ (28)
 where $t_{0ac}$ is chosen to be the instant in which accretion starts, $r_{0ac}$ is the apparent horizon at time $t=t_{0ac}$, $r_{b}$ is the geometric locus of the black hole apparent horizon, and $M_{i}$ is defined as $M_{i}=M_{MS}^{\left(0\right)}\left(t_{0ac},r_{0ac}\right)+M_{ac}\left(t_{0ac},r_{0ac}\right)$, with $M_{MS}^{\left(0\right)}\left(t,r\right)$ given by Eq.
@@
 A solution similar to Eq.
-(LABEL:M_solution) appears in Refs.
-[babichev2012backreaction, campos2025dynamical].
+(28) appears in Refs.
+[42, 43].
 However, in the present context, we are dealing with a slightly more general scenario, since both the accretion term and the energy density are allowed to depend on time.
 Moreover, it is worth highlighting that working with the McVittie metric allows us to obtain solutions analogous to that of Refs.
-[babichev2012backreaction, campos2025dynamical], but expressed in a physical time coordinate rather than in an Eddington–Finkelstein–type time coordinate.
+[42, 43], but expressed in a physical time coordinate rather than in an Eddington–Finkelstein–type time coordinate.
 The expression for $\mathcal{M}_{MS}$ reveals the contribution of the accretion of matter, described by $\mathcal{T}_{\nu}^{\text{ }\mu}$, to the initial effective mass $M_{i}$.
-The first integral in (LABEL:M_solution) corresponds to the contribution from the evolution of the accreting matter flux, evaluated in the region bounded by $r_{0ac}$ over the interval $t-t_{0ac}$.
+The first integral in (28) corresponds to the contribution from the evolution of the accreting matter flux, evaluated in the region bounded by $r_{0ac}$ over the interval $t-t_{0ac}$.
 The second integral represents the time-dependent contribution from the energy density contained between the initial apparent horizon before accretion, $r_{0ac}$ , and the apparent horizon displaced due to accretion, $r_{b}$.
-With the solution (LABEL:M_solution) at hand, we can now effectively compute the effects of accretion.
-We only need to evaluate the integrals in Eq.(LABEL:M_solution) in terms of the energy–momentum tensors.
-We shall assume that the tensors $\mathcal{T}_{\nu}^{\ \mu}$ satisfying Einstein’s equations (15) correspond to a perfect fluid, given by
+With the solution (28) at hand, we can now effectively compute the effects of accretion.
+We only need to evaluate the integrals in Eq.
+(28) in terms of the energy–momentum tensors.
+We shall assume that the tensors $\mathcal{T}_{\nu}^{\ \mu}$ satisfying Einstein’s equations (19)-(22) correspond to a perfect fluid, given by
 $\mathcal{T}_{\nu}^{\ \mu}=\left({T}_{\nu}^{\ \mu}\right)^{(0)}+\left(\rho_{ac}+p_{ac}\right)u_{\nu}u^{\mu}+p_{ac}\delta_{\nu}^{\mu},$ (29)
@@
 To address this, we shall employ the perturbative approach used in Ref.
-[babichev2012backreaction].
+[42].
 We assume that the effects of the accreting fluid are much smaller than those of the background fluid, so that the energy–momentum tensor of the accreting fluid can be evaluated in the background metric $g_{\mu\nu}^{\left(0\right)}$ given by (1).
-Consequently, the tensor $\mathcal{T}_{\nu}^{\text{ }\mu}$ and the solution (LABEL:M_solution) will be considered as a first-order perturbation, namely
+Consequently, the tensor $\mathcal{T}_{\nu}^{\text{ }\mu}$ and the solution (28) will be considered as a first-order perturbation, namely
 $\mathcal{T}_{\nu}^{\text{ }\mu}\rightarrow\left(T_{\nu}^{\text{ }\mu}\right)^{\left(1\right)},$ (32)
@@
 Eqs. (30) and (31) then reduce to:
-$\left(T_{0}^{\text{ }0}\right)^{\left(1\right)}\left(t,r_{b}\right)$ $=$ $-\frac{3H_{0}^{2}\left(t\right)}{8\pi}+\frac{\left(\rho_{ac}+p_{ac}\right)H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}{v-2H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}$ (28)
-$+p_{ac}$
+$\left(T_{0}^{\text{ }0}\right)^{\left(1\right)}\left(t,r_{b}\right)=-\frac{3H_{0}^{2}\left(t\right)}{8\pi}+\frac{\left(\rho_{ac}+p_{ac}\right)H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}{v-2H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}+p_{ac}$ (35)
 and
 $\left(T_{0}^{\text{ }1}\right)^{\left(1\right)}\left(t,r_{b}\right)=\frac{\left(\rho_{ac}+p_{ac}\right)H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}v}{v-2H_{0}^{2}\left(t\right)\left(r_{b}^{\left(0\right)}\left(t\right)\right)^{2}}.$ (36)
-Under this perturbative approximation, Eq. (LABEL:M_solution) expressed in terms of the Eqs. (28) and (29) is exact at the initial time $t=t_{0ac}$, such that
+Under this perturbative approximation, Eq. (28) expressed in terms of the Eqs. (35) and (36) is exact at the initial time $t=t_{0ac}$, such that
 $g_{00}\left(t_{0ac},r_{0ac}\right)=g_{00}^{\left(0\right)}\left(t_{0ac},r_{b}^{\left(0\right)}\left(t_{0ac}\right)\right),$ (37)
 and provides an approximation for $t>t_{0ac}$.
-We may then analyze the integrals appearing in Eq.
-(LABEL:M_solution) using the definitions from Eqs.
-(19) and (20).
-We are going to define the first integral in Eq.
-(LABEL:M_solution), after substituting Eq.
-(29), as
+We may then analyze the integrals appearing in Eq. (28) using the definitions from Eqs. (26) and (27). We are going to define the first integral in Eq. (28), after substituting Eq. (36), as
 $F\left(t\right)\equiv\int_{t_{0ac}}^{t}\mathcal{A}\left(\bar{t},r_{0ac}\right)~d\bar{t}.$ (38)
 Since our goal is to study the accretion of the GCG, before evaluating the second integral in Eq.
-(LABEL:M_solution), we anticipate that $\rho_{ac}$ and $p_{ac}$ with which we will work depend only on time.
+(28), we anticipate that $\rho_{ac}$ and $p_{ac}$ with which we will work depend only on time.
 Evidently, the same assumption holds for the radial velocity of the fluid, $v$, once our goal is to evaluate the accretion at a specific point, the black hole horizon, i.e., $v(t,r)=v(t,r^{(1)}_{b})$.
@@
 (27) becomes
-$\int_{r_{0ac}}^{r_{b}^{\left(1\right)}}\mathcal{B}\left(t,\bar{r}\right)~d\bar{r}$ $=$ $-4\pi\left(T_{0}^{\text{ }0}\left(t\right)\right)^{\left(1\right)}\int_{r_{0ac}}^{r_{b}^{\left(1\right)}}\bar{r}^{2}~d\bar{r}=$ (32)
-$=$ $\frac{H_{1}^{2}\left(t\right)}{2}\left(r_{b}^{\left(1\right)}\right)^{3}+G\left(t\right),$
+$\int_{r_{0ac}}^{r_{b}^{\left(1\right)}}\mathcal{B}\left(t,\bar{r}\right)~d\bar{r}=-4\pi\left(T_{0}^{\text{ }0}\left(t\right)\right)^{\left(1\right)}\int_{r_{0ac}}^{r_{b}^{\left(1\right)}}\bar{r}^{2}~d\bar{r}=\frac{H_{1}^{2}\left(t\right)}{2}\left(r_{b}^{\left(1\right)}\right)^{3}+G\left(t\right),$ (39)
 where we define
@@
 $G\left(t\right)\equiv-\frac{H_{1}^{2}\left(t\right)}{2}r_{0ac}^{3}.$ (42)
-Finally, the solution given in Eq. (LABEL:M_solution) can be expressed as
+Finally, the solution given in Eq. (28) can be expressed as
 $M_{MS}^{\left(1\right)}\left(t,r\right)=M_{1}\left(t\right)+\frac{H_{1}^{2}\left(t\right)}{2}\left(r_{b}^{(1)}\right)^{3},$ (43)
@@
 $H_{1}^{2}\left(t\right)\left(r_{b}^{\left(1\right)}\right)^{3}-r_{b}^{\left(1\right)}+2M_{1}\left(t\right)=0.$ (46)
-This equation is structurally analogous to the cubic Eq. (11). Its physically meaningful solutions, in agreement with (12a) and (12), are given by
+This equation is structurally analogous to the cubic Eq. (13). Its physically meaningful solutions, in agreement with (14) and (2), are given by
 $r_{b}^{\left(1\right)}\left(t\right)$ $=\frac{2}{\sqrt{3}H_{1}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right),$
-$r_{c}^{\left(1\right)}\left(t\right)$ $=$ $\frac{1}{H_{1}\left(t\right)}\cos\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right)$
-$-\frac{1}{\sqrt{3}H_{1}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right)$
-. (39b)
+$r_{c}^{\left(1\right)}\left(t\right)$ $=\frac{1}{H_{1}\left(t\right)}\cos\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right)$ (48)
+$-\frac{1}{\sqrt{3}H_{1}\left(t\right)}\sin\left(\frac{\arcsin\left(3\sqrt{3}M_{1}\left(t\right)H_{1}\left(t\right)\right)}{3}\right).$
 Therefore, within the adopted perturbative framework, we obtain analytical expressions for the temporal evolution, expressed in terms of a physical time coordinate, of both the black hole and cosmological apparent horizons as a result of the accretion process.
 Moreover, the conditions required for the solutions in Eqs.
-(39) to be real are similar to those discussed in the previous section for Eqs.
-(12a) and (12), now for $M_{0}\rightarrow M_{1}\left(t\right)$ e $H_{0}\left(t\right)\rightarrow H_{1}\left(t\right)$.
+(LABEL:r1b)-(48) to be real are similar to those discussed in the previous section for Eqs.
+(14) and (2), now for $M_{0}\rightarrow M_{1}\left(t\right)$ and $H_{0}\left(t\right)\rightarrow H_{1}\left(t\right)$.
 Naturally, the analytical character of the solutions $r_{b}^{\left(1\right)}\left(t\right)$ and $r_{c}^{\left(1\right)}\left(t\right)$ depends on whether $M_{1}\left(t\right)$ and $H_{1}\left(t\right)$ can also be treated analytically.
 Therefore, we shall investigate the explicit solutions for specific $\rho_{ac}$ and $p_{ac}$.
-IV Applications
-IV.1 Dark energy accretion
+4 Applications
+4.1 Dark energy accretion
 Before analyzing the GCG, let us first consider a simpler case: the accretion of a fluid whose equation of state corresponds to that of dark energy, namely,
@@
 A similar equation appears in Ref.
-[babichev2012backreaction], but in the case where $H_{0}=0$.
+[42], but in the case where $H_{0}=0$.
 In Ref.
-[babichev2012backreaction] treatment, Einstein’s equations are evaluated in a vacuum background, that is, $G_{\mu\gamma}\left[g_{\mu\gamma}^{(0)}\right]=0$.
+[42] treatment, Einstein’s equations are evaluated in a vacuum background, that is, $G_{\mu\gamma}\left[g_{\mu\gamma}^{(0)}\right]=0$.
 In contrast, our zeroth-order solution to Einstein’s equations is given by the McVittie equations.
-IV.2 Generalized Chaplygin Gas
+4.2 Generalized Chaplygin Gas
 Let us now analyze the accretion of the GCG, whose equation of state is given by Eq.
@@
 $\rho^{(dS)}_{GCG}\approx A^{1/\left(1+\alpha\right)}\Rightarrow p^{(dS)}_{GCG}=-\rho^{(dS)}_{GCG}$. (55)
-Furthermore, there exists an effective equation of state in the intermediate regime between the matter-dominated and de Sitter phases, which can be obtained by expanding Eq.
-(7) as in Ref.
-[bento2002generalized]:
-$\rho_{GCG}$ $\approx$ $A^{1/\left(1+\alpha\right)}+\left(\frac{1}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)}+$...,
-$p_{GCG}$ $\approx$ $-A^{1/\left(1+\alpha\right)}+\left(\frac{\alpha}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)}+$...
-Eqs. (47) are essential, since within the perturbative framework we are using, the contribution of the accreting fluid must be much smaller than those of the background fluid.
+Furthermore, there exists an effective equation of state in the intermediate regime between the matter-dominated and de Sitter phases, which can be obtained by expanding Eq. (8) as in Ref. [10]:
+$\rho_{GCG}$ $\approx A^{1/\left(1+\alpha\right)}+\left(\frac{1}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)}+$..., (56)
+$p_{GCG}$ $\approx-A^{1/\left(1+\alpha\right)}+\left(\frac{\alpha}{1+\alpha}\right)\frac{B}{A^{\alpha/\left(1+\alpha\right)}}a^{-3\left(1+\alpha\right)}+$... (57)
+Eqs. (56)-(57) are essential, since within the perturbative framework we are using, the contribution of the accreting fluid must be much smaller than those of the background fluid.
 However, since we are assuming that the GCG drives the cosmic expansion, thus determining the background metric, while we are also interested in analyzing its accretion, we shall adopt the following strategy: we associate the dominant term in Eq.
-(47) with the background fluid characterized by $H_{0}$, and the remaining term with the perturbation.
+(56)-(57) with the background fluid characterized by $H_{0}$, and the remaining term with the perturbation.
 For example, in a de Sitter universe:
@@
 We shall now analyze each case separately.
-IV.2.1 Matter dominated era
+4.2.1 Matter-dominated era
 In a universe governed by the GCG during its matter-dominated phase, from Eq. (54) we have
 $H_{0}\left(t\right)=\frac{2}{3t}.$ (60)
-According to our perturbative construction, we shall assume that the accreting fluid obeys the subdominant term of the equation of state given in Eq. (47). Thus, its equation of state reads
+According to our perturbative construction, we shall assume that the accreting fluid obeys the subdominant term of the equation of state given in Eq. (56)-(57). Thus, its equation of state reads
 $\rho_{ac}=-p_{ac}=A^{1/\left(1+\alpha\right)}.$ (61)
 Consequently, from Eqs. (38), (41), and (44), we obtain
-$F\left(t\right)$ $=$ $0,$ (52a)
-$H_{1}^{2}\left(t\right)$ $=$ $\frac{4}{9t^{2}}+\frac{8\pi A^{1/\left(1+\alpha\right)}}{3},$ (52b)
-$M_{1}\left(t\right)$ $=$ $M_{0}+M_{ac}+\left(H_{0}^{2}\left(t_{0ac}\right)-H_{1}^{2}\left(t\right)\right)\frac{r_{0ac}^{3}}{2},$
+$F\left(t\right)$ $=0,$ (62)
+$H_{1}^{2}\left(t\right)$ $=\frac{4}{9t^{2}}+\frac{8\pi A^{1/\left(1+\alpha\right)}}{3},$ (63)
+$M_{1}\left(t\right)$ $=M_{0}+M_{ac}+\left(H_{0}^{2}\left(t_{0ac}\right)-H_{1}^{2}\left(t\right)\right)\frac{r_{0ac}^{3}}{2},$ (64)
 where, for an energy density given by (61), $M_{ac}=4\pi r_{ac}^{3}\rho_{ac}/3$. Therefore, the effective mass expressed as a first-order perturbation as given by Eq. (43), becomes
@@
 To this end, we compare Eqs.
-(12a) and (12) with Eqs.
-(39).
-From Eqs. (12a) and (12), as discussed in section II, when the equation
+(14) and (2) with Eqs.
+(LABEL:r1b)-(48).
+From Eqs. (14) and (2), as discussed in section 2, when the equation
 $3\sqrt{3}M_{0}H_{0}\left(t\right)=1$ (66)
@@
 Similarly to the non-accreting case, from Eqs.
-(39) it follows that the horizons are formed at instant $t_{0ac}$ satisfying
+(LABEL:r1b)-(48) it follows that the horizons are formed at instant $t_{0ac}$ satisfying
 $3\sqrt{3}M_{1}\left(t_{0ac}\right)H_{1}\left(t_{0ac}\right)=1.$ (69)
-Using Eqs. (52b) and (LABEL:M1_matter), Eq. (57) admits, as its positive solution,
+Using Eqs. (63) and (64), Eq. (69) admits, as its positive solution,
 $t_{0ac}=\frac{t_{0}}{\sqrt{1-72\pi M_{0}^{2}A^{1/\left(1+\alpha\right)}}}=\frac{t_{0}}{\sqrt{1-72\pi M_{0}^{2}\rho_{ac}}},$ (70)
@@
 Figure 1: Apparent black hole and cosmological horizon formation time as a function of AA (solid lines), as given by (70).
-The dashed line represent the instant of formation of the apparent horizons without available matter for accretion, $t_{0}$, given by 56.
+The dashed line represent the instant of formation of the apparent horizons without available matter for accretion, $t_{0}$, given by 68.
 The mass parameter $M_{0}$ is set to 11.
@@
 The horizons obtained from Eqs.
-(39), together with Eqs.
-(52b) and (LABEL:M1_matter), are shown in Fig.
+(LABEL:r1b)-(48), together with Eqs.
+(63) and (64), are shown in Fig.
 2 for different values of $A$.
@@
 After a sufficiently long time, such that $t>>t_{0ac}$, the horizons given by Eqs.
-(12a) and (12) behave as
-$r_{b}^{\left(0\right)}\left(t\right)\approx 2M_{0}\text{, \ }r_{c}^{\left(0\right)}\left(t\right)\approx\frac{1}{H_{0}}-M_{0}\approx\frac{3t}{2}-M_{0}.$ (59)
-Similarly, for Eqs. (39), we find that for time scales such that $t>>t_{0ac}$, the behavior of the horizons in the presence of accretion is given by
-$r_{b}^{\left(1\right)}\left(t\right)$ $\approx$ $2M_{1}\left(t\right)\approx 3M_{0}-36\pi A^{1/(1+\alpha)}M_{0}^{3},$ (60a)
-$r_{c}^{\left(1\right)}\left(t\right)$ $\approx$ $\frac{1}{H_{1}\left(t\right)}-M_{1}\left(t\right)\approx\left(\frac{4}{9t^{2}}+\frac{4\pi A^{1/(1+\alpha)}}{3}\right)^{-1/2}$ (60b)
-$-\frac{3M_{0}}{2}+18\pi A^{1/(1+\alpha)}M_{0}^{3},$
+(14) and (2) behave as
+$r_{b}^{\left(0\right)}\left(t\right)$ $\approx 2M_{0}$ (71)
+$r_{c}^{\left(0\right)}\left(t\right)$ $\approx\frac{1}{H_{0}}-M_{0}\approx\frac{3t}{2}-M_{0}.$ (72)
+Similarly, for Eqs. (LABEL:r1b)-(48), we find that for time scales such that $t>>t_{0ac}$, the behavior of the horizons in the presence of accretion is given by
+$r_{b}^{\left(1\right)}\left(t\right)$ $\approx 2M_{1}\left(t\right)\approx 3M_{0}-36\pi A^{1/(1+\alpha)}M_{0}^{3},$ (73)
+$r_{c}^{\left(1\right)}\left(t\right)$ $\approx\frac{1}{H_{1}\left(t\right)}-M_{1}\left(t\right)\approx\left(\frac{4}{9t^{2}}+\frac{4\pi A^{1/(1+\alpha)}}{3}\right)^{-1/2}-\frac{3M_{0}}{2}$ (74)
+$+18\pi A^{1/(1+\alpha)}M_{0}^{3},$
 where we used the fact that $r_{0ac}\equiv r_{b}^{(1)}(t_{0ac})=3M_{0}$.
 For the black hole horizon with accretion to be larger than in the case without accretion, the condition $A<\left(12\pi M_{0}^{2}\right)^{-\left(1+\alpha\right)}$ must be satisfied, according to Eq.
-(60a).
+(73).
 However, due to the constraint previously imposed by Eq.
@@
 Observationally consistent models typically give $10^{-60}{$\mathrm{m}$}^{-2(1+\alpha)}<A<10^{-54}{$\mathrm{m}$}^{-2(1+\alpha)}$ for $\alpha\leq 0.1$, as in Refs.
-[aurich2018concordance, barreiro2008wmap].
+[12, 13].
 Using enhanced values in our simulations makes the accretion effects numerically visible while preserving the qualitative behavior of the horizon evolution.
@@
 The inflow of accreting matter locally increases the energy density and consequently reduces the rate at which cosmic expansion decelerates, as shown in Eqs.
-(60a) and (52b).
+(73) and (63).
 The resulting higher effective expansion rate counteracts the growth of the local gravitational term, namely, the apparent black hole horizon, as expressed in Eqs.
-(60a) and (LABEL:M1_matter), thereby delaying the instant at which the apparent horizon condition is satisfied.
-IV.2.2 de Sitter era
+(73) and (64), thereby delaying the instant at which the apparent horizon condition is satisfied.
+4.2.2 de Sitter era
 In a de Sitter–type universe governed by the GCG, the Hubble parameter is given by
@@
 According to our perturbative approach, since the background is determined by an equation of state of the form $\rho^{(dS)}_{GCG}=-p^{(dS)}_{GCG}=A^{1/\left(1+\alpha\right)}$, we assume that the accreting fluid corresponds to the subdominant term in Eq.
-(47), namely
+(56)-(57), namely
 $p_{ac}=\alpha\rho_{ac}=\left(\frac{\alpha}{1+\alpha}\right)\frac{B}{A^{\alpha/(1+\alpha)}}a^{-3\left(1+\alpha\right)}.$ (78)
@@
 As discussed previously, Refs.
-[aurich2018concordance, barreiro2008wmap] typically give $10^{-60}{$\mathrm{m}$}^{-2(1+\alpha)}<A<10^{-54}{$\mathrm{m}$}^{-2(1+\alpha)}$ for $\alpha\leq 0.1$.
+[12, 13] typically give $10^{-60}{$\mathrm{m}$}^{-2(1+\alpha)}<A<10^{-54}{$\mathrm{m}$}^{-2(1+\alpha)}$ for $\alpha\leq 0.1$.
 If we conservatively assume an initial accretion velocity $v_{0}=0.01c_{s}$, where $c_{s}=\sqrt{{\partial p}/{\partial\rho}}$ is the sound speed derived from the equation of state (78), we obtain $v_{0}=0.01\alpha^{1/2}$.
@@
 Therefore, from Eq. (38) we get
-$F\left(t\right)$ $=$ $64\pi H_{A}^{2}M_{0}^{4}(1+\alpha)\int_{t_{0ac}}^{t}\rho_{ac}d\bar{t}=$ (66)
-$=$ $\frac{64\pi H_{A}M_{0}^{4}}{3}\left(\rho_{ac0}-\rho_{ac}\left(t\right)\right)$
+$F\left(t\right)=64\pi H_{A}^{2}M_{0}^{4}(1+\alpha)\int_{t_{0ac}}^{t}\rho_{ac}d\bar{t}=\frac{64\pi H_{A}M_{0}^{4}}{3}\left(\rho_{ac0}-\rho_{ac}\left(t\right)\right)$ (80)
 where $\rho_{ac0}$ denotes the density given by Eq. (78) at the instant $t_{0ac}$ when accretion begins.
@@
 Finally, for the function $M_{1}\left(t\right)$ defined in Eq. (44), we obtain
-$M_{1}\left(t\right)$ $=$ $M_{0}+\frac{32\pi M^{3}_{0}}{3}\rho_{ac0}+18\pi\alpha M_{0}^{3}\rho_{ac}$ (71)
-$+\frac{64H_{A}M_{0}^{4}\pi}{3}\left(\rho_{ac0}-\rho_{ac}\left(t\right)\right),$
+$M_{1}\left(t\right)=M_{0}+\frac{32\pi M^{3}_{0}}{3}\rho_{ac0}+18\pi\alpha M_{0}^{3}\rho_{ac}+\frac{64H_{A}M_{0}^{4}\pi}{3}\left(\rho_{ac0}-\rho_{ac}\left(t\right)\right),$ (85)
 where we used $M_{ac}(t_{0ac})=4\pi r^{3}_{0ac}\rho_{ac0}/3$.
@@
 (78), the black hole horizon becomes larger in the presence of accretion than in its absence, while the cosmological horizon becomes correspondingly smaller.
-V Conclusions
+5 Conclusions
 In this work, we investigated the accretion of cosmic dark fluids, particularly the GCG, onto cosmologically coupled black holes described by the McVittie metric.
@@
 This complementary behavior highlights the subtle interplay between local and global dynamics in cosmologically coupled systems.
+The formalism developed here is sufficiently general to be extended to a variety of related scenarios.
+In particular, a more detailed investigation of the backreaction effects on the fluid dynamics itself would be of considerable interest, especially concerning the structure of critical sonic points and the transonic nature of the flow in a cosmological background [28, 29, 44].
+Moreover, the analysis may be generalized to other perfect fluids coupled to the expansion of the universe through alternative equations of state.
+Such extensions would allow one to assess the robustness of the qualitative features reported here.
 Overall, since the McVittie metric provides a bridge between cosmological expansion and local gravitational collapse, our formalism could be particularly relevant for the study of primordial black holes, whose formation and evolution are naturally embedded in an expanding universe.
 Understanding how accretion and backreaction operate in such early-universe scenarios could shed light on the role of cosmologically coupled black holes in structure formation and in the broader dynamics of the dark sector.
-Acknowledgements
+Acknowledgments
+The work of L.F.R.
+is supported by FCT (Fundação para a Ciência e Tecnologia, Portugal) through the grant 2024.06413.BD.
 The work of O.B.
-is partially supported by FCT (Fundação para a Ciência e Tecnologia, Portugal) through the project 2024.00252.CERN.
-The work of L.F.R.
-is supported by FCT through the grant 2024.06413.BD.
+is partially supported by FCT through the project 2024.00252.CERN.
 Both authors are partially supported by UID/04650 - Centro de Física das Universidades do Minho e do Porto.